The mean of marks in Statistics of 100 students of a class was 72. The mean of marks of boys was 75 , while their number was 70. Find out the mean marks of girls in the class.



1
Expert's answer
2022-03-07T17:02:07-0500

There are 100-70=30 girls in the class.

Let x be the mean marks of the girls.

"\frac{70*75+30*x}{100}=72."

"30x=100*72-70*75=1950."

"x=\frac{1950}{30}=65."